### v1

# Nevai’s condition for measures with unbounded supports

Abstract.
We study Nevai’s condition from the theory of orthogonal polynomials on the real line. We prove that a large class of measures with unbounded Jacobi parameters satisfies Nevai’s condition locally uniformly on the support of the measure away from a finite explicit set. This allows us to give applications to relative uniform and weak asymptotics of Christoffel–Darboux kernel on the diagonal and to limit theorems of unconventionally normalized global linear statistics of orthogonal polynomial ensembles.
Key words and phrases:
Orthogonal polynomials; Christoffel–Darboux kernel; linear statistics of orthogonal polynomial ensembles
2020 Mathematics Subject Classification:
Primary 42C05; Secondary 47B36
1. Introduction
Let $\mu$ be a Borel probability measure on the real line with infinite support such that all the moments
$\int_{\mathbb{R}}x^{n}{\>\rm d}\mu(x),\quad n\geq 0$
exist and are finite. Let $L^{2}(\mu)$ be the usual Hilbert space associated with the scalar product
$\langle f,g\rangle_{L^{2}(\mu)}=\int_{\mathbb{R}}f(x)\overline{g(x)}{\>\rm d}\mu(x).$
By performing Gram–Schmidt orthonormalization process on the sequence $(x^{n}:n\geq 0)$ of monomials we get the sequence of orthonormal polynomials $(p_{n}:n\geq 0)$. By the construction, for any $n\geq 0$ the polynomial $p_{n}$ is of exact degree $n$, its leading coefficient is positive, and for any $n\geq 0,m\geq 0$ we have
(1.1) $\langle p_{n},p_{m}\rangle_{L^{2}(\mu)}=\begin{cases}1&n=m,\\ 0&\text{otherwise}.\end{cases}$
Since the polynomial $xp_{n}$ is of degree $n+1$, by orthogonality we easily get the following three-term recurrence relation
(1.2) $p_{0}(x)\equiv 1,\quad p_{1}(x)=\frac{x-b_{0}}{a_{0}},$
(1.3) $xp_{n}(x)=a_{n-1}p_{n-1}(x)+b_{n}p_{n}(x)+a_{n}p_{n+1}(x),\quad n\geq 1,$
where
$a_{n}=\langle xp_{n},p_{n+1}\rangle_{L^{2}(\mu)}>0,\quad b_{n}=\langle xp_{n},p_{n}\rangle_{L^{2}(\mu)}\in\mathbb{R},\quad n\geq 0$
are the so-called Jacobi parameters of $\mu$.
One can reverse this procedure and for any given sequences $(a_{n}:n\geq 0),(b_{n}:n\geq 0)$ of positive and real numbers, respectively, define the sequence of polynomials $(p_{n}:n\geq 0)$ by the formulas (1.2) and (1.3). Then Favard’s theorem (see e.g. [25, Remark 5.15]) asserts existence of a probability measure $\mu$ on the Borel subsets of the real line such that the sequence $(p_{n}:n\geq 0)$ satisfies the formula (1.1). The measure $\mu$ need not to be unique. If it is unique, the measure $\mu$ is called determinate. Otherwise it is indeterminate. If Jacobi parameters are bounded, i.e.
(1.4) $\sup_{n\geq 0}(a_{n}+|b_{n}|)<\infty,$
then $\mu$ is determinate and $\operatornamewithlimits{supp}(\mu)$ is compact. More generally, Carleman’s condition
(1.5) $\sum_{n=0}^{\infty}\frac{1}{a_{n}}=\infty$
implies determinacy of $\mu$.
A central object in the theory of orthogonal polynomials are the Christoffel–Darboux kernels defined for any $n\geq 1$ by
(1.6) $K_{n}(x,y;\mu)=\sum_{j=0}^{n-1}p_{j}(x)p_{j}(y),\quad x,y\in\mathbb{R}.$
We shall omit $\mu$ in the notation if the measure $\mu$ is clear from the context. The kernels $K_{n}$ are reproducing in $L^{2}(\mu)$ for $\mathbb{P}_{n-1}$, the subspace of all complex polynomials with degree less or equal to $n-1$. It means that
(1.7) $p(x)=\int_{\mathbb{R}}K_{n}(x,y)p(y){\>\rm d}\mu(y),\qquad p\in\mathbb{P}_{n-1},\ x\in\mathbb{R}.$
They can be expressed in terms of the polynomials $p_{n-1}$ and $p_{n}$ through the Christoffel–Darboux formula
$K_{n}(x,y)=a_{n-1}\begin{cases}\frac{p_{n}(x)p_{n-1}(y)-p_{n-1}(x)p_{n}(y)}{x-y}&\text{if }x\neq y,\\ p_{n}^{\prime}(x)p_{n-1}(x)-p_{n-1}^{\prime}(x)p_{n}(x)&\text{otherwise}.\end{cases}$
For some applications of Christoffel–Darboux kernel in orthogonal polynomials, spectral theory, random matrix theory and data analysis, see e.g. [21], [27], [16] and [14], respectively.
Nevai in his monograph [20] considered in detail properties of orthogonal polynomials with Jacobi parameters satisfying
(1.8) $\lim_{n\to\infty}a_{n}=\frac{1}{2}\quad\text{and}\quad\lim_{n\to\infty}b_{n}=0.$
He proved in [20, Theorem 4.2.14] that under (1.8) the following weak convergence of probability measures
[…]
$\lim_{n\to\infty}\int_{\mathbb{R}}f{\>\rm d}\omega_{n}=\int_{\mathbb{R}}f{\>\rm d}\omega_{\infty},\quad f\in\mathcal{C}_{b}(\mathbb{R}).$
Here by $\mathcal{C}_{b}(\mathbb{R})$ we denote the Banach space of continuous bounded functions on the real line equipped with the supremum norm $\|\cdot\|_{\infty}$.
The formula (1.9) allows to approximate the measure $\mu$ provided one understands the pointwise behaviour of $K_{n}(x,x)$. Along these lines, consider the following decomposition of the measure $\mu$
(1.10) ${\>\rm d}\mu={\mu^{\prime}_{\mathrm{ac}}}{\>\rm d}\lambda+{\>\rm d}{\mu_{\mathrm{s}}},$
where the measure ${\mu_{\mathrm{s}}}$ is singular with respect to the Lebesgue measure $\lambda$. Suppose that $\operatornamewithlimits{supp}(\mu)=[-1,1]$ and ${\mu^{\prime}_{\mathrm{ac}}}(x)>0$ for $\lambda$-a.a. $x\in[-1,1]$. Then by Rakhmanov’s theorem (see [24, Theorem 2] and also [6]) the formula (1.8) holds true. Moreover, by [17, Theorem 12.1] it holds
(1.11) $\lim_{n\to\infty}\int_{-1}^{1}\bigg{|}\frac{1}{n}K_{n}(x,x){\mu^{\prime}_{\mathrm{ac}}}(x)-\frac{1}{\pi\sqrt{1-x^{2}}}\bigg{|}{\>\rm d}\lambda(x)=0$
(cf. (1.9)). Furthermore, if a stronger Szegő’s condition is satisfied, i.e.
(1.12) $\int_{-1}^{1}\frac{\log{\mu^{\prime}_{\mathrm{ac}}}(x)}{\sqrt{1-x^{2}}}{\>\rm d}\lambda(x)>-\infty,$
then one even has
(1.13) $\lim_{n\to\infty}\frac{1}{n}K_{n}(x,x)=\frac{1}{{\mu^{\prime}_{\mathrm{ac}}}(x)}\frac{1}{\pi\sqrt{1-x^{2}}},\quad\text{$\lambda$-a.a. }x\in(-1,1)$
(see [18, Theorem 5]). Let us mention that generalizations of (1.9), (1.11) and (1.13) to compactly supported measures $\mu$ satisfying some regularity conditions have been accomplished in [30], [28] and [42], respectively.
Since the condition (1.12) is quite restrictive and far from being necessary for (1.13) to hold, for a well-understood measure $\mu$ it is natural to consider new measures of the form
(1.14) ${\>\rm d}\mu_{g}(x):=g(x){\>\rm d}\mu(x),$
where $g$ is a positive $\mu$-measurable function such that the measure $\mu_{g}$ has all moments finite. By normalizing, one can assume that $\mu_{g}$ is a probability measure. If the behaviour of $K_{n}(x,x;\mu)$ is known, it seems to be useful to study the expression
(1.15) $\frac{K_{n}(x,x;\mu_{g})}{K_{n}(x,x;\mu)}.$
To understand the behaviour of (1.15) Nevai in [20, Section 6.2] considered for any nonnegative $\mu$-measurable function $f$ and any $n\geq 1$ a new function $G_{n}[f]:\mathbb{R}\to[0,\infty]$ defined by
(1.16) $G_{n}[f](x)=\frac{1}{K_{n}(x,x)}\int_{\mathbb{R}}K_{n}^{2}(x,y)f(y){\>\rm d}\mu(y),\quad x\in\mathbb{R}.$
According to Proposition 2.1 the right-hand side of the formula (1.16) is a bounded rational function for any $f\in\mathcal{L}^{\infty}(\mu)$, so we shall extend the definition (1.16) to such functions. Let us recall that the space $\mathcal{L}^{\infty}(\mu)$ consists of $\mu$-essentially bounded measurable functions with the essential supremum seminorm $\|\cdot\|_{\mathcal{L}^{\infty}(\mu)}$.
Let us recall central notions of [4].
Definition 1.1.
Let $\mu$ be a probability measure on the real line with infinite support having all moments finite. We say that the measure $\mu$ satisfies Nevai’s condition at some $x\in\mathbb{R}$ if
(1.17) $\lim_{n\to\infty}G_{n}[f](x)=f(x),\quad f\in\mathcal{C}_{b}(\mathbb{R}).$
If $K$ is a bounded subset of $\mathbb{R}$ and
(1.18) $\lim_{n\to\infty}\sup_{x\in K}\big{|}G_{n}[f](x)-f(x)\big{|}=0,\quad f\in\mathcal{C}_{b}(\mathbb{R}),$
then we say that the measure $\mu$ satisfies Nevai’s condition uniformly on $K$.
The conditions (1.17) and (1.18) have been introduced by Nevai in [20, Section 6.2]. His remarkable observation (see Corollary 3.1) states that if $g:\mathbb{R}\to(0,\infty)$ is such that $g,1/g\in\mathcal{L}^{\infty}(\mu)$ and $g$ is continuous at some $x\in\mathbb{R}$, then
(1.19) $\lim_{n\to\infty}\frac{K_{n}(x,x;\mu_{g})}{K_{n}(x,x;\mu)}=\frac{1}{g(x)}$
provided (1.17) is satisfied. Moreover, if additionally $g$ is continuous on some compact $K\subset\mathbb{R}$ and (1.18) holds, then the convergence (1.19) is uniform on $K$.
Nevai’s conditions in the case of compactly supported measures $\mu$ have been studied in detail. It turns out that for such measures the condition (1.17) is equivalent to the subexponential growth of the sequence $(p_{n}(x):n\geq 0)$, which means
(1.20) $\lim_{n\to\infty}\frac{p_{n}^{2}(x)}{K_{n}(x,x)}=0,$
see [4, Theorem 1.2]. The condition (1.20) allowed to prove that Nevai’s condition holds uniformly on $\operatornamewithlimits{supp}(\mu)$ for a large class of compactly supported measures, see e.g. [4, Theorem 8.2]. In the same paper there were also presented examples of compactly supported measures such that Nevai’s condition fails for some $x\in\operatornamewithlimits{supp}(\mu)$. In the article [15] for any measure $\mu$ with compact support and any $f\in\mathcal{L}^{\infty}(\lambda)$ it was proven that
$\lim_{n\to\infty}\lambda\big{(}\{x\in\mathbb{R}:{\mu^{\prime}_{\mathrm{ac}}}(x)>0,|G_{n}[f](x)-f(x)|>\epsilon\}\big{)}=0,\quad\epsilon>0$
and for any $g:\mathbb{R}\to(0,\infty)$ satisfying $g,1/g\in\mathcal{L}^{\infty}(\mu)$
$\lim_{n\to\infty}\lambda\bigg{(}\bigg{\{}x\in\mathbb{R}:{\mu^{\prime}_{\mathrm{ac}}}(x)>0,\bigg{|}\frac{K_{n}(x,x;\mu_{g})}{K_{n}(x,x;\mu)}-\frac{1}{g(x)}\bigg{|}>\epsilon\bigg{\}}\bigg{)}=0,\quad\epsilon>0.$
Let us also mention the article [3], where Nevai’s condition found applications for limit theorems of linear statistics of a class of determinantal point processes, called orthogonal polynomial ensembles.
The present article is concerned on the case of measures with unbounded supports. On the page 251 of the article [4] it was suggested that study Nevai’s condition in this case seems to be an interesting challenge. It was also mentioned that their methods completely break down for the measures from the article [5]. It turns out that for these measures Nevai’s condition holds locally uniformly on $\operatornamewithlimits{supp}(\mu)\setminus\{0\}$ (see Section 4.1.4). Next, [3, Theorem 4.5] allows to prove Nevai’s condition for arbitrary measures at Lebesgue points of $\mu$ under the additional sine kernel universality of the Christoffel–Darboux kernels. In view of the recent breakthrough in sine kernel universality (see [7]) this additional hypothesis follows from the previous one.
In Section 2 we present sufficient conditions for uniform Nevai’s condition. In particular, Theorem 2.4 is a uniform variant of [3, Theorem 4.5]. Theorem 2.3 shows that Nevai’s condition always holds on the support of discrete part of any determinate measure. Interestingly, the same implication holds also for some (the so-called $N$-extremal) indeterminate measures.
Section 3 is devoted to some implications of uniform Nevai’s condition for $\mu$ combined with some information on the behaviour of $(K_{n}(x,x;\mu):n\geq 1)$. We study there a number of properties related to the measure $\mu_{g}$ defined in (1.14) under the additional assumption: $g,1/g\in\mathcal{C}_{b}(\mathbb{R})$. In particular, in Section 3.1 we relate pointwise, uniform or weak asymptotics of $(K_{n}(x,x;\mu_{g}):n\geq 1)$ to the same kind of asymptotics of $(K_{n}(x,x;\mu):n\geq 1)$. Finally, in Section 3.2 we study limit theorems for unconventionally normalized global statistics of orthogonal polynomial ensembles.
In the last Section 4 we show examples of measures $\mu$ with unbounded supports defined in terms of Jacobi parameters which satisfies Nevai’s condition locally uniformly on $\operatornamewithlimits{supp}(\mu)\setminus X$, where $X$ is a finite exceptional set. In particular, we cover a large class of periodically modulated and blended Jacobi parameters. Since for such Jacobi parameters we know the asymptotic behaviour of Christoffel–Darboux kernels, this allows us to apply the tools from Section 3 to measures of the form (1.14). This seems to be an important step forward, since in general no understanding of Jacobi parameters of the measures $\mu_{g}$ are known.
Acknowledgements
I would like to thank Bartosz Trojan for useful discussions on the topic of the present article.
2. Properties of Nevai’s operators
Let us start with a useful consequence of the reproducing property (1.7) of the Christoffel–Darboux kernels. Namely, since $K_{n}(x,\cdot)$ is a polynomial of degree $n-1$ with respect to the second variable, we get by (1.7)
(2.1) $K_{n}(x,x)=\int_{\mathbb{R}}K_{n}^{2}(x,y){\>\rm d}\mu(y),\qquad x\in\mathbb{R}.$
The following proposition implies that the assignment $\mathcal{L}^{\infty}(\mu)\ni f\mapsto G_{n}[f]$ defines a contractive operator $G_{n}:\mathcal{L}^{\infty}(\mu)\to\mathcal{C}_{b}(\mathbb{R})$ (cf. [20, Properties 6.2.1(iii)]). Therefore, it seems natural to call $(G_{n}:n\geq 1)$ as Nevai’s operators.
Proposition 2.1.
Let $f\in\mathcal{L}^{\infty}(\mu)$. Then $G_{n}[f]$ is a rational function on $\mathbb{R}$ satisfying
(2.2) $\|G_{n}[f]\|_{\infty}\leq\|f\|_{\mathcal{L}^{\infty}(\mu)}.$
Proof.
Let $f\in\mathcal{L}^{\infty}(\mu)$. Then by Hölder’s inequality together with (2.1) gives
(2.3) $\big{|}G_{n}[f](x)\big{|}\leq\int_{\mathbb{R}}\frac{K_{n}^{2}(x,y)}{K_{n}(x,x)}|f(y)|{\>\rm d}\mu(y)\leq\|f\|_{\mathcal{L}^{\infty}(\mu)},\quad x\in\mathbb{R},$
which shows that $G_{n}[f]$ is a well-defined function. Next, by (1.6)
(2.4) $G_{n}[f](x)=\frac{1}{K_{n}(x,x)}\sum_{0\leq j,j\leq n-1}p_{j}(x)p_{j^{\prime}}(x)\int_{\mathbb{R}}p_{j}(y)p_{j^{\prime}}(y)f(y){\>\rm d}\mu(y).$
By Hölder’s inequality applied twice
$\bigg{|}\int_{\mathbb{R}}p_{j}(y)p_{j^{\prime}}(y)f(y){\>\rm d}\mu(y)\bigg{|}$ $\leq\|f\|_{\mathcal{L}^{\infty}(\mu)}\int_{\mathbb{R}}|p_{j}(y)p_{j^{\prime}}(y)|{\>\rm d}\mu(y)$
$\leq\|f\|_{\mathcal{L}^{\infty}(\mu)}\|p_{j}\|_{L^{2}(\mu)}\|p_{j^{\prime}}\|_{L^{2}(\mu)}$
$=\|f\|_{\mathcal{L}^{\infty}(\mu)},$
where in the last equality we used (1.1). Thus, the expression (2.4) is a well-defined rational function. Since its denominator do not vanish, it is a continuous function and (2.3) implies (2.2). The proof is complete. ∎
2.1. Sufficient conditions for Nevai’s condition
It turns out that Nevai’s condition has an alternative definition in the measure theoretical language. Namely, for a given $x\in\mathbb{R}$ let us define the sequence of measures
(2.5) ${\>\rm d}\omega^{x}_{n}(y)=\frac{K_{n}^{2}(x,y)}{K_{n}(x,x)}{\>\rm d}\mu(y).$
By (2.1) these measures are probability measures. Moreover, by (1.16) and (2.5) one has
(2.6) $G_{n}[f](x)=\int_{\mathbb{R}}f(y){\>\rm d}\omega^{x}_{n}(y).$
[…]
We start with the following lemma which is a part of the proof of [3, Theorem 4.4].
Lemma 2.2.
Let $K$ be a bounded subset of $\mathbb{R}$. If there exists $\eta_{0}>0$ such that for each $\eta\in(0,\eta_{0}]$,
(2.7) $\liminf_{n\to\infty}\inf_{x\in K}\omega^{x}_{n}\big{(}[x-\eta,x+\eta]\big{)}=1,$
then $\mu$ satisfies Nevai’s condition uniformly on $K$.
Proof.
Let $f\in\mathcal{C}_{b}(\mathbb{R})$. By (2.6) we need to show that
(2.8) $\lim_{n\to\infty}\sup_{x\in K}\bigg{|}\int_{\mathbb{R}}\big{(}f(y)-f(x)\big{)}{\>\rm d}\omega_{n}^{x}(y)\bigg{|}=0.$
To see this, we write
$\bigg{|}\int_{\mathbb{R}}\big{(}f(y)-f(x)\big{)}{\>\rm d}\omega_{n}^{x}(y)\bigg{|}$ $\leq\int_{\mathbb{R}\setminus[x-\eta,x+\eta]}|f(y)-f(x)|{\>\rm d}\omega_{n}^{x}(y)$
$\phantom{\leq}+\int_{[x-\eta,x+\eta]}|f(y)-f(x)|{\>\rm d}\omega_{n}^{x}(y).$
For the first integral we have
$\sup_{x\in K}\int_{\mathbb{R}\setminus[x-\eta,x+\eta]}|f(y)-f(x)|{\>\rm d}\omega_{n}^{x}(y)$ $\leq 2\|f\|_{\infty}\sup_{x\in K}\Big{(}1-\nu_{n}^{x}\big{(}[x-\eta,x+\eta]\big{)}\Big{)}$
(2.9) $=2\|f\|_{\infty}\Big{(}1-\inf_{x\in K}\nu_{n}^{x}\big{(}[x-\eta,x+\eta]\big{)}\Big{)}.$
The second integral can be estimated as follows
(2.10) $\sup_{x\in K}\int_{[x-\eta,x+\eta]}|f(y)-f(x)|{\>\rm d}\omega_{n}^{x}(y)\leq\sup_{x\in K}\sup_{h\in[-\eta,\eta]}|f(x+h)-f(x)|.$
Now, in view of (2.7), the right-hand side of (2.9) tends to 00 for any $\eta\in(0,\eta_{0}]$. Next, the right-hand side of (2.10) tends to 00 as $\eta\downarrow 0$ by the uniform continuity of $f$ on the compact set( [^1] For any $X\subset\mathbb{R}$ by $\operatorname{cl}({X})$ we denote its closure in Euclidean topology.) $\tilde{K}:=\operatorname{cl}({K})+[-\delta,\delta]$. Hence, (2.8) follows. ∎
The following theorem states that Nevai’s condition is always satisfied at each point of the support of the discrete part of the measure $\mu$.
Theorem 2.3.
Let $\mu$ be a probability measure on the real line with all moments finite. Suppose that the corresponding sequence $(p_{n}:n\geq 0)$ of orthogonal polynomials forms an orthonormal basis in $L^{2}(\mu)$ (in particular it is satisfied when the measure $\mu$ is determinate). If for some $x\in\mathbb{R}$ it holds $\mu(\{x\})>0$, then the measure $\mu$ satisfies Nevai’s condition at $x$.
Proof.
For every $\eta>0$ we have
$\omega^{x}_{n}\big{(}[x-\eta,x+\eta]\big{)}\geq\frac{K_{n}^{2}(x,x;\mu)}{K_{n}(x,x;\mu)}\mu(\{x\})=K_{n}(x,x;\mu)\mu(\{x\}).$
According to M. Riesz theorem (see e.g. [26, Theorem 2.14]) the orthogonal polynomials $(p_{n}:n\geq 0)$ form an orthonormal basis in $L^{2}(\mu)$ iff the measure $\mu$ is either determinate or the so-called $N$-extremal. Then
$\lim_{n\to\infty}K_{n}(x,x;\mu)=\frac{1}{\mu(\{x\})},$
(see e.g. [26, Corollary 2.6] in the determinate case and [26, Theorem 2.13] in the N-extremal case) and the result follows from Lemma 2.2 applied to $K=\{x\}$. ∎
By adapting the proofs of [3, Theorems 4.4 and 4.5], we obtain the following theorem.
Theorem 2.4.
For a probability measure $\mu$ on the real line having all moments finite consider the decomposition (1.10) and define
(2.11) $\tilde{K}_{n}(x,y)=\sqrt{{\mu^{\prime}_{\mathrm{ac}}}(x){\mu^{\prime}_{\mathrm{ac}}}(y)}K_{n}(x,y;\mu),\quad x,y\in\mathbb{R}.$
Suppose that the measure $\mu$ is absolutely continuous on some open $U\subset\mathbb{R}$. If for some bounded $K\subset U$ we have
(2.12) $\lim_{n\to\infty}\inf_{x\in K}\tilde{K}_{n}(x,x)=\infty,$
and for any $x\in K$
(2.13) $\liminf_{n\to\infty}\frac{1}{\tilde{K}^{2}_{n}(x,x)}\tilde{K}_{n}^{2}\bigg{(}x,x+\frac{t}{\tilde{K}_{n}(x,x)}\bigg{)}\geq\operatorname{sinc}^{2}(\pi t),\quad\text{$\lambda$-a.a. }t\in\mathbb{R},$
then $\mu$ satisfies Nevai’s condition uniformly on $K$. Moreover, the measure $\mu$ is determinate.
Proof.
The condition (2.12) implies that the measure $\mu$ is determinate (see e.g. [25, Theorem 6.16(v)]). We are going to prove that the hypotheses of Lemma 2.2 are satisfied on $K$. Let $\delta_{0}>0$ be such that
$\bigcup_{x\in K}[x-\delta_{0},x+\delta_{0}]\subset U$
To prove (2.7), we observe that for $\delta\in(0,\delta_{0}]$ by (2.11)
$\omega^{x}_{n}\big{(}[x-\delta,x+\delta]\big{)}$ $=\int_{x-\delta}^{x+\delta}\frac{K_{n}^{2}(x,y)}{K_{n}(x,x)}{\mu^{\prime}_{\mathrm{ac}}}(y){\>\rm d}y$
$=\int_{x-\delta}^{x+\delta}\frac{\tilde{K}_{n}^{2}(x,y)}{\tilde{K}_{n}(x,x)}{\>\rm d}y$
$=\int_{-\delta\tilde{K}_{n}(x,x)}^{\delta\tilde{K}_{n}(x,x)}\frac{\tilde{K}_{n}^{2}\Big{(}x,x+\frac{t}{\tilde{K}_{n}(x,x)}\Big{)}}{\tilde{K}^{2}_{n}(x,x)}{\>\rm d}t.$
By (2.12) for any $M>0$ there exists $N_{0}>0$ such that for any $n\geq N_{0}$ we have $\inf_{x\in K}\tilde{K}_{n}(x,x)\geq M/\delta$. Thus, by Fatou’s Lemma and (2.13)
$\liminf_{n\to\infty}\inf_{x\in K}\omega^{x}_{n}\big{(}[x-\delta,x+\delta]\big{)}$ $\geq\liminf_{n\to\infty}\int_{-M}^{M}\frac{\tilde{K}_{n}^{2}\Big{(}x,x+\frac{t}{\tilde{K}_{n}(x,x)}\Big{)}}{\tilde{K}^{2}_{n}(x,x)}{\>\rm d}t$
$\geq\int_{-M}^{M}\liminf_{n\to\infty}\frac{\tilde{K}_{n}^{2}\Big{(}x,x+\frac{t}{\tilde{K}_{n}(x,x)}\Big{)}}{\tilde{K}^{2}_{n}(x,x)}{\>\rm d}t$
$\geq\int_{-M}^{M}\operatorname{sinc}^{2}(\pi t){\>\rm d}t.$
Since $M>0$ was arbitrary and
$\int_{-\infty}^{\infty}\operatorname{sinc}^{2}(\pi t){\>\rm d}t=1$
we get (2.7) what we needed to show. ∎
Remark 2.5.
Let us recall that according to [7, Theorem 1.2] and their Remark (b) there is a choice of ${\mu^{\prime}_{\mathrm{ac}}}$ such that for any $x$ being a Lebesgue point of $\mu$ such that ${\mu^{\prime}_{\mathrm{ac}}}(x)>0$ (in particular for ${\mu_{\mathrm{ac}}}$-a.a. $x\in\mathbb{R}$) the limit
(2.14) $\lim_{n\to\infty}\frac{1}{\tilde{K}_{n}(x,x)}\tilde{K}_{n}\Big{(}x+\frac{u}{\tilde{K}_{n}(x,x)},x+\frac{v}{\tilde{K}_{n}(x,x)}\Big{)}=\operatorname{sinc}\big{(}\pi(u-v)\big{)}$
holds locally uniformly with respect to $u,v\in\mathbb{R}^{2}$. It implies that for such $x$ the condition (2.13) holds.
Remark 2.6.
Suppose that $K\subset\mathbb{R}$ is compact. The condition (2.12) holds if for some positive sequence $(\rho_{n}:n\in\mathbb{N})$ tending to infinity and some function $\upsilon:K\to(0,\infty)$
(2.15) $\lim_{n\to\infty}\frac{1}{\rho_{n}}\tilde{K}_{n}(x,x)=\upsilon(x)$
uniformly on $K$.
3. Applications of Nevai’s condition
In this Section we shall study implications of uniform Nevai’s condition for $\mu$ combined with some information on the behaviour of $(K_{n}(x,x;\mu):n\geq 1)$.
3.1. Comparative asymptotics of Christoffel–Darboux kernels on the diagonal
Let $\mu$ be a probability measure on $\mathbb{R}$ with infinite support and all moments finite. Let $g:\mathbb{R}\to(0,\infty)$ be such that $g,1/g\in\mathcal{L}^{\infty}(\mu)$. Then by [20, p. 76, Theorem 6.2.3] we have
(3.1) $\frac{1}{G_{n}[g](x)}\leq\frac{K_{n}(x,x;g{\>\rm d}\mu)}{K_{n}(x,x;\mu)}\leq G_{n}\bigg{[}\frac{1}{g}\bigg{]}(x).$
Thus, by (2.2) we obtain
(3.2) $\frac{1}{\|g\|_{\mathcal{L}^{\infty}(\mu)}}\leq\frac{K_{n}(x,x;g{\>\rm d}\mu)}{K_{n}(x,x;\mu)}\leq\bigg{\|}\frac{1}{g}\bigg{\|}_{\mathcal{L}^{\infty}(\mu)}.$
Corollary 3.1.
Suppose that $\mu$ satisfies Nevai’s condition at some $x\in\mathbb{R}$. Let $g:\mathbb{R}\to(0,\infty)$ be such that $g,1/g\in\mathcal{L}^{\infty}(\mu)$. If $g$ is continuous at $x$, then
(3.3) $\lim_{n\to\infty}\frac{K_{n}(x,x;g{\>\rm d}\mu)}{K_{n}(x,x;\mu)}=\frac{1}{g(x)}$
Moreover, if $\mu$ satisfies uniform Nevai’s condition on some compact $K\subset\mathbb{R}$ and $g$ is continuous on $K$, then the convergence in (3.3) is uniform on $K$.
Proof.
It follows from applying (1.17) (or (1.18)) to (3.1). ∎
Corollary 3.1 can be used to obtain the following.
Theorem 3.2.
Suppose that for some positive sequence $(\rho_{n}:n\geq 1)$ tending to infinity, there is a positive Radon measure $\nu$ satisfying $\int_{\mathbb{R}}\frac{1}{1+x^{2}}{\>\rm d}\nu(x)<\infty$ such that
[…]
Suppose that there is $X\subset\mathbb{R}$ with $\nu(X)=0$ such that $\mu$ satisfies Nevai’s condition locally uniformly around each $x\in\mathbb{R}\setminus X$. Let $g:\mathbb{R}\to(0,\infty)$ be such that $g,1/g\in\mathcal{C}_{b}(\mathbb{R})$. Then
[…]
Proof.
Let $f\in\mathcal{C}_{b}(\mathbb{R})$. We can write
(3.6) $\frac{1}{\rho_{n}}\int_{\mathbb{R}}f(x)\frac{K_{n}(x,x;g{\>\rm d}\mu)}{1+x^{2}}g(x){\>\rm d}\mu(x)=\int_{\mathbb{R}}h_{n}(x){\>\rm d}\eta_{n}(x),$
where
$h_{n}(x)=f(x)g(x)\frac{K_{n}(x,x;g{\>\rm d}\mu)}{K_{n}(x,x;\mu)},\quad{\>\rm d}\eta_{n}=\frac{1}{\rho_{n}}\frac{K_{n}(x,x;\mu)}{1+x^{2}}{\>\rm d}\mu(x).$
[…]
$\|h_{n}\|_{\infty}\leq\|f\|_{\infty}\|g\|_{\infty}\bigg{\|}\frac{1}{g}\bigg{\|}_{\infty}<\infty$
the sequence $(h_{n}:n\geq 1)$ is uniformly integrable with respect to $(\eta_{n}:n\geq 1)$ (cf. [8, Definition 2.4]). By Corollary 3.1
$\lim_{n\to\infty}h_{n}(x)=f(x)$
locally uniformly with respect to $x\in\mathbb{R}\setminus X$. Thus
$\lim_{\begin{subarray}{c}n\to\infty\\ x^{\prime}\to x\end{subarray}}h_{n}(x)=f(x),\quad x\in\mathbb{R}\setminus X.$
Therefore, the hypotheses of [8, Corollary 5.1] are satisfied, and consequently,
$\lim_{n\to\infty}\int_{\mathbb{R}}h_{n}(x){\>\rm d}\eta_{n}(x)=\int_{\mathbb{R}}f(x)\frac{1}{1+x^{2}}{\>\rm d}\nu(x),$
what in view (3.6) implies (3.5) and ends the proof. ∎
3.2. Law of large numbers for global linear statistics of orthogonal polynomial ensembles
In this section we shall study limit theorems for linear statistics of an important subclass of determinantal point processes, which contains many models considered in physics, statistical mechanics, probability theory and combinatorics, see the survey [13] for more details.
An orthogonal polynomial ensemble of size $n\in\mathbb{N}$ is a probability measure $\mathbb{P}^{(n)}$ on $\mathbb{R}^{n}$ of the form
(3.7) ${\>\rm d}\mathbb{P}^{(n)}(\lambda_{1},\ldots,\lambda_{n})=\frac{1}{Z_{n}}\prod_{1\leq i<j\leq n}(\lambda_{j}-\lambda_{i})^{2}{\>\rm d}\mu(\lambda_{1})\ldots{\>\rm d}\mu(\lambda_{n}),$
where $\mu$ is a probability measure on $\mathbb{R}$ with all moments finite and $Z_{n}>0$ is a normalizing constant. It can be shown that (3.7) can be rewritten to
(3.8) ${\>\rm d}\mathbb{P}^{(n)}(\lambda_{1},\ldots,\lambda_{n})=\frac{1}{n!}\det\big{[}K_{n}(\lambda_{i},\lambda_{j};\mu)\big{]}_{i,j=1}^{n}{\>\rm d}\mu(\lambda_{1})\ldots{\>\rm d}\mu(\lambda_{n}),$
see e.g. [13, Lemma 2.8].
Suppose that a random vector $\big{\{}\lambda_{i}^{(n)}\big{\}}_{i=1}^{n}$ has the distribution $\mathbb{P}^{(n)}$. One can associate to it a random measure
$\Xi^{(n)}=\sum_{i=1}^{n}\delta_{\lambda_{i}^{(n)}}.$
For any $\mu$-measurable function $f:\mathbb{R}\to\mathbb{R}$ one considers a linear statistic of $\Xi^{(n)}$ defined by
$\Xi^{(n)}f=\sum_{i=1}^{n}f\big{(}\lambda_{i}^{(n)}\big{)}.$
From the determinantal structure of (3.8) we have
(3.9) $\mathbb{E}[\Xi^{(n)}f]=\int_{\mathbb{R}}f(x)K_{n}(x,x){\>\rm d}\mu(x)$
and
(3.10) $\operatorname{Var}[\Xi^{(n)}f]=\int_{\mathbb{R}}f^{2}(x)K_{n}(x,x){\>\rm d}\mu(x)-\int_{\mathbb{R}}\int_{\mathbb{R}}f(x)f(y)K_{n}^{2}(x,y){\>\rm d}\mu(x){\>\rm d}\mu(y),$
see e.g. [3, Appendix A.2].
In what follows, we shall assume that the random measures $(\Xi^{(n)}:n\geq 1)$ are defined on the same probability space $(\Omega,\mathcal{F},\mathbb{P})$.
Lemma 3.3.
Suppose that for some positive sequence $(\rho_{n}:n\geq 1)$ tending to infinity, there is a positive Radon measure $\nu$ satisfying $\int_{\mathbb{R}}\frac{1}{1+x^{2}}{\>\rm d}\nu(x)<\infty$ such that
[…]
Then for any $f\in\mathcal{C}_{b}(\mathbb{R})$ satisfying $\sup_{x\in\mathbb{R}}(1+x^{2})|f(x)|<\infty$ we have
(3.12) $\lim_{n\to\infty}\frac{1}{\rho_{n}}\mathbb{E}[\Xi^{(n)}f]=\int_{\mathbb{R}}f{\>\rm d}\nu,$
(3.13) $\limsup_{n\to\infty}\frac{1}{\rho_{n}}\operatorname{Var}[\Xi^{(n)}f]<\infty.$
In particular, for any $\epsilon>0$
(3.14) $\lim_{n\to\infty}\mathbb{P}\bigg{(}\bigg{|}\frac{1}{\rho_{n}}\Xi^{(n)}f-\int_{\mathbb{R}}f(x){\>\rm d}\nu(x)\bigg{|}\geq\epsilon\bigg{)}=0.$
Proof.
The formula (3.12) follows immediately from (3.9) and (3.11). Next, in view of (1.16)
(3.15) $\int_{\mathbb{R}}\int_{\mathbb{R}}f(x)f(y)K_{n}^{2}(x,y){\>\rm d}\mu(x){\>\rm d}\mu(y)=\int_{\mathbb{R}}f(x)G_{n}[f](x)K_{n}(x,x){\>\rm d}\mu(x).$
Since (2.2) we have
(3.16) $\bigg{|}\int_{\mathbb{R}}f(x)G_{n}[f](x)K_{n}(x,x){\>\rm d}\mu(x)\bigg{|}\leq\lVert{f}\rVert_{\infty}\int_{\mathbb{R}}|f(x)|K_{n}(x,x){\>\rm d}\mu(x)$
Therefore, by (3.10) and (3.11) we get
$\limsup_{n\to\infty}\frac{1}{\rho_{n}}\operatorname{Var}[\Xi^{(n)}f]\leq\int_{\mathbb{R}}f^{2}{\>\rm d}\nu+\lVert{f}\rVert_{\infty}\int_{\mathbb{R}}|f|{\>\rm d}\nu<\infty,$
which proves (3.13). Finally, by Chebyshev inequality for any $\epsilon>0$
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{\rho_{n}}\Xi^{(n)}f-\mathbb{E}\bigg{[}\frac{1}{\rho_{n}}\Xi^{(n)}f\bigg{]}\bigg{|}\geq\epsilon\bigg{)}$ $\leq\frac{1}{\epsilon^{2}}\operatorname{Var}\bigg{[}\frac{1}{\rho_{n}}\Xi^{(n)}f\bigg{]}$
(3.17) $\leq\frac{1}{\epsilon^{2}}\frac{1}{\rho_{n}^{2}}\operatorname{Var}[\Xi^{(n)}f].$
Thus, the formula (3.14) follows from (3.13) and (3.12). ∎
Under additional hypotheses we can prove that the limit (3.13) is equal to 00.
Lemma 3.4.
Suppose that the hypotheses of Lemma 3.3 are satisfied. Let $f\in\mathcal{C}_{b}(\mathbb{R})$ be such that $\sup_{x\in\mathbb{R}}(1+x^{2})|f(x)|<\infty$. If
(3.18) $\liminf_{\begin{subarray}{c}n\to\infty\\ x^{\prime}\to x\end{subarray}}\big{(}f(x^{\prime})G_{n}[f](x^{\prime})\big{)}\geq f^{2}(x),\quad\nu\text{-a.a. }x\in\mathbb{R},$
where $G_{n}[f]$ is defined in (1.16), then
(3.19) $\lim_{n\to\infty}\frac{1}{\rho_{n}}\operatorname{Var}[\Xi^{(n)}f]=0.$
Proof.
By (3.10) and (3.15) we can write
(3.20) $\frac{1}{\rho_{n}}\operatorname{Var}[\Xi^{(n)}f]=\frac{1}{\rho_{n}}\int_{\mathbb{R}}f^{2}(x)K_{n}(x,x){\>\rm d}\mu-\int_{\mathbb{R}}h_{n}(x){\>\rm d}\eta_{n}(x)$
where
(3.21) $h_{n}(x)=(1+x^{2})f(x)G_{n}[f](x),\quad{\>\rm d}\eta_{n}=\frac{1}{\rho_{n}}\frac{K_{n}(x,x)}{1+x^{2}}{\>\rm d}\mu(x)$
Since $\operatorname{Var}[\Xi^{(n)}f]\geq 0$ by taking the limit superior and using (3.12) we get
$\limsup_{n\to\infty}\int_{\mathbb{R}}h_{n}(x){\>\rm d}\eta_{n}(x)\leq\int_{\mathbb{R}}f^{2}{\>\rm d}\nu.$
Thus, (3.19) will follow once we prove
(3.22) $\liminf_{n\to\infty}\int_{\mathbb{R}}h_{n}(x){\>\rm d}\eta_{n}(x)\geq\int_{\mathbb{R}}f^{2}{\>\rm d}\nu.$
[…]
$\lVert{h_{n}}\rVert_{\infty}\leq\lVert{f}\rVert_{\infty}\cdot\sup_{x\in\mathbb{R}}(1+x^{2})|f(x)|<\infty,$
the sequence $(h_{n}:n\geq 1)$ is uniformly integrable with respect to $(\eta_{n}:n\geq 1)$ (cf. [8, Definition 2.4]). Therefore, the hypotheses of [8, Theorem 2.2] are satisfied. Therefore,
$\liminf_{n\to\infty}\int_{\mathbb{R}}h_{n}(x){\>\rm d}\eta_{n}(x)\geq\int_{\mathbb{R}}\liminf_{\begin{subarray}{c}n\to\infty\\ x^{\prime}\to x\end{subarray}}h_{n}(x^{\prime})\frac{1}{1+x^{2}}{\>\rm d}\nu(x).$
Now, by using (3.18) and (3.21) we get (3.22), what concludes the proof. ∎
Remark 3.5.
Let $f\in\mathcal{C}_{b}(\mathbb{R})$. If the measure $\mu$ satisfies the Nevai’s condition uniformly in some closed interval $K\subset\mathbb{R}$, then the condition
$\lim_{\begin{subarray}{c}n\to\infty\\ x^{\prime}\to x\end{subarray}}G_{n}[f](x^{\prime})=f(x)$
holds for any $x\in\operatorname{int}(K)$. To see this, it is enough to notice that for any $x,x^{\prime}\in K$
$\big{|}G_{n}[f](x^{\prime})-f(x)\big{|}\leq\sup_{x^{\prime}\in K}\big{|}G_{n}[f](x^{\prime})-f(x^{\prime})\big{|}+\big{|}f(x^{\prime})-f(x)\big{|}$
we have by (1.18)
$\lim_{\begin{subarray}{c}n\to\infty\\ x^{\prime}\to x\end{subarray}}\big{|}G_{n}[f](x^{\prime})-f(x)\big{|}\leq\lim_{x^{\prime}\to x}|f(x^{\prime})-f(x)|=0,$
by continuity of $f$ at $x$.
Once the limit (3.13) is equal to zero, we sometimes can improve the convergence (3.14) to the $\mathbb{P}$-almost sure convergence. For a related result, see [9, Corollary 1.4]
Theorem 3.6.
Under the hypotheses of Lemma 3.4 we have that for any $\epsilon>0$ there exists $M\geq 1$ such that for $n\geq M$
(3.23) $\mathbb{P}\bigg{(}\bigg{|}\frac{1}{\rho_{n}}\Xi^{(n)}f-\frac{1}{\rho_{n}}\mathbb{E}[\Xi^{(n)}f]\bigg{|}\geq\epsilon\bigg{)}\leq 2\exp\bigg{(}-\frac{\epsilon}{6\lVert{f}\rVert_{\infty}}\rho_{n}\bigg{)}.$
In particular,
(3.24) $\mathbb{P}\bigg{(}\lim_{n\to\infty}\frac{1}{\rho_{n}}\Xi^{(n)}f=\int_{\mathbb{R}}f{\>\rm d}\nu\bigg{)}=1$
provided
[…]
Proof.
Since
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{\rho_{n}}\Xi^{(n)}f-\frac{1}{\rho_{n}}\mathbb{E}[\Xi^{(n)}f]\bigg{|}\geq\epsilon\bigg{)}=\mathbb{P}\big{(}\big{|}\Xi^{(n)}f-\mathbb{E}[\Xi^{(n)}f]\big{|}\geq\epsilon\rho_{n}\big{)}.$
We have by [3, Theorem 2.1] the bound (3.23) for all $n\geq 1$ such that
(3.26) $\epsilon\geq\frac{2A}{3\lVert{f}\rVert_{\infty}}\frac{1}{\rho_{n}}\operatorname{Var}[\Xi^{(n)}f],$
where $A>0$ is some constant. In view of (3.19) the right-hand side of (3.26) tends to 00. So there is a constant $M\geq 1$ such that the inequality (3.26) is satisfied for all $n\geq M$, which implies (3.23). Next, if (3.25) is satisfied, then by Borel-Cantelli lemma and (3.12) we get (3.24). ∎
Remark 3.7.
Suppose that $\operatornamewithlimits{supp}(\mu)=[-1,1]$ and ${\mu^{\prime}_{\mathrm{ac}}}(x)>0$ for a.a. $x\in(-1,1)$. If $f\in\mathcal{L}^{\infty}(\lambda)$, then by [22, Theorem 2]
$\lim_{n\to\infty}\frac{1}{n}\operatorname{Var}[\Xi^{(n)}f]=0.$
As it was mentioned in the introduction, by Rakhmanov’s theorem one has (1.8), so the formula (1.9) holds true. Thus, if additionally the set of discontinuity points of $f$ has Lebesgue measure zero, then by repeating the argument of the proof of Theorem 3.6 we get
$\mathbb{P}\bigg{(}\lim_{n\to\infty}\frac{1}{n}\Xi^{(n)}f=\int_{-1}^{1}f(x)\frac{1}{\pi\sqrt{1-x^{2}}}{\>\rm d}\lambda(x)\bigg{)}=1.$
Remark 3.8.
Suppose that the sequence $(\rho_{n}:n\in\mathbb{N})$ is strictly increasing and sublinear, i.e. the sequence $r_{n}=\rho_{n+1}-\rho_{n}$ is convergent to 00. Then Raabe’s test seems to be a convenient sufficient condition for (3.25) to hold. Indeed, notice that
$\lim_{n\to\infty}\frac{1}{r_{n}}\bigg{(}\frac{\textrm{e}^{-c\rho_{n}}}{\textrm{e}^{-c\rho_{n+1}}}-1\bigg{)}=\lim_{n\to\infty}\frac{\textrm{e}^{cr_{n}}-1}{r_{n}}=c>0.$
Thus,
$\lim_{n\to\infty}n\bigg{(}\frac{\textrm{e}^{-c\rho_{n}}}{\textrm{e}^{-c\rho_{n+1}}}-1\bigg{)}=c\lim_{n\to\infty}nr_{n}.$
Hence, by Raabe’s test, the condition (3.25) is satisfied provided $\lim_{n\to\infty}nr_{n}=\infty$.
4. Examples
In this section we shall consider unbounded Jacobi parameters defined in terms of periodic sequences. It turns out that it will be useful to put some regularity conditions. More specifically, we say that a sequence $(x_{n}:n\geq 1)$ belongs to $\mathcal{D}_{r}$ for some integer $r\geq 1$, if it is bounded and for each $j\in\{1,\ldots,r\}$
$\sum_{n=1}^{\infty}|\Delta^{j}x_{n}|^{\tfrac{r}{j}}<\infty$
where
$\Delta^{0}x_{n}$ $=x_{n}$
$\Delta^{j}x_{n}$ $=\Delta^{j-1}x_{n+1}-\Delta^{j-1}x_{n},\quad j\geq 1.$
This class has been introduced in [31]. Let us recall that $\mathcal{D}_{1}$ consists of sequences of bounded variation and for any $r\geq 1$ one has $\mathcal{D}_{r}\subsetneq\mathcal{D}_{r+1}$. Given an integer $N\geq 1$ we say that $(x_{n}:n\geq 1)\in\mathcal{D}_{r}^{N}$ if for any $i=0,1,\ldots,N-1$ we have $(x_{nN+i}:n\geq 1)\in\mathcal{D}_{r}$.
It will be useful to rewrite the recurrence relation (1.3) to the form
(4.1) $\begin{pmatrix}p_{n}(x)\\ p_{n+1}(x)\end{pmatrix}=B_{n}(x)\begin{pmatrix}p_{n-1}(x)\\ p_{n}(x)\end{pmatrix},\quad\text{where}\quad B_{n}(x)=\begin{pmatrix}0&1\\ -\frac{a_{n-1}}{a_{n}}&\frac{x-b_{n}}{a_{n}}\end{pmatrix},\quad n\geq 1.$
The matrix $B_{n}(x)$ is usually ($1$-step) transfer matrix.
By $(\alpha_{n}:n\in\mathbb{Z}),(\beta_{n}:n\in\mathbb{Z})$ we denote $N$-periodic sequences of positive and real numbers, respectively. Moreover, we define
(4.2) $\mathfrak{B}_{n}(x)=\begin{pmatrix}0&1\\ -\frac{\alpha_{n-1}}{\alpha_{n}}&\frac{x-\beta_{n}}{\alpha_{n}}\end{pmatrix},\quad n\in\mathbb{Z}.$
4.1. Periodic modulations
We say that Jacobi parameters $(a_{n}),(b_{n})$ are $N$-periodically modulated, if there exist $N$-periodic sequences $(\alpha_{n}),(\beta_{n})$ of positive and real numbers, respectively such that
(4.3) $\lim_{n\to\infty}\bigg{|}\frac{a_{n-1}}{a_{n}}-\frac{\alpha_{n-1}}{\alpha_{n}}\bigg{|}=0,\quad\lim_{n\to\infty}\bigg{|}\frac{b_{n}}{a_{n}}-\frac{\beta_{n}}{\alpha_{n}}\bigg{|}=0,\quad\lim_{n\to\infty}a_{n}=\infty.$
Next, we define $N$-step transfer matrix by
$X_{n}(x)=B_{N+n-1}(x)\ldots B_{n+1}(x)B_{n}(x),\quad n\geq 1.$
Notice that (4.3) implies that
$\lim_{j\to\infty}X_{jN+i}(x)=\mathfrak{X}_{i}(0),\quad i=0,1,\ldots,N-1,$
where
(4.4) $\mathfrak{X}_{n}(x)=\mathfrak{B}_{N+n-1}(x)\ldots\mathfrak{B}_{n+1}(x)\mathfrak{B}_{n}(x),\quad n\in\mathbb{Z}.$
and $\mathfrak{B}_{n}(x)$ is defined in (4.2). This class has been introduced in [10] and it is studied ever since (see e.g. [38, 37, 39, 40, 36, 41, 34, 32, 33, 35, 29, 5, 11, 19, 23, 2] and the references therein). Notice that by (4.3) the condition (1.4) is violated. Therefore, measures $\mu$ corresponding to $N$-periodically modulated Jacobi parameters have unbounded supports. It turns out that, under some regularity assumptions on Jacobi parameters, a few of fundamental properties of $\mu$ depend crucially on the matrix $\mathfrak{X}_{0}(0)$. More specifically, one can distinguish the four cases:
  1. (I)
    if $\operatorname{tr}\mathfrak{X}_{0}(0)\in(-2,2)$, then the measure $\mu$ is purely absolutely continuous on $\mathbb{R}$ with positive continuous density;
  2. (II)
    if $\operatorname{tr}\mathfrak{X}_{0}(0)\in\{-2,2\}$, then we have two subcases:
    1. (a)
      if $\mathfrak{X}_{0}(0)$ is diagonalizable, then there is a compact interval $K\subset\mathbb{R}$ such that the measure $\mu$ is purely absolutely continuous on $\mathbb{R}\setminus K$ with positive continuous density and it is purely discrete in the interior of $K$;
    2. (b)
      if $\mathfrak{X}_{0}(0)$ is not diagonalizable, then there is a possibly degenerated infinite open interval $\Lambda\subset\mathbb{R}$ such that the measure $\mu$ is purely absolutely continuous on $\Lambda$ with positive continuous density and purely discrete on $\mathbb{R}\setminus\Lambda$;
  3. (III)
    if $\operatorname{tr}\mathfrak{X}_{0}(0)\in\mathbb{R}\setminus[-2,2]$, then the measure $\mu$ is purely discrete with the support having no finite accumulation points.
One can describe these four cases geometrically. Specifically, we have
$(\operatorname{tr}\mathfrak{X}_{0})^{-1}\big{(}(-2,2)\big{)}=\bigcup_{j=1}^{N}I_{j},$
where $I_{j}$ are disjoint open non-empty bounded intervals whose closures might touch each other. Let $I_{j}=(x_{2j-1},x_{2j})$, where the sequence $(x_{k}:k=1,2,\ldots,2N)$ is increasing. Then we are in the case I if 00 belongs to some interval $I_{j}$, in the case IIII(a) if 00 lies on the boundary of exactly two intervals, in the case IIII(b) if 00 lies on the boundary of exactly one interval and in the case III in the remaining cases. An example for $N=4$ is presented in Figure 1.
$x_{4}=x_{5}$$x_{1}$$x_{2}$$x_{3}$$x_{6}$$x_{7}$$x_{8}$
Figure 1. An example for $N=4$. If $0=x_{4}$, then we are in the case IIII(a), while $0\in\{x_{1},x_{2},x_{3},x_{6},x_{7},x_{8}\}$ corresponds to the case IIII(b).
Below, we shall show that, under some regularity conditions, measures $\mu$ corresponding to $N$-periodically modulated Jacobi parameters satisfy Nevai’s condition locally uniformly on $\operatornamewithlimits{supp}(\mu)\setminus X$, where $X$ is some explicit finite set. This show that results from Section 3 can be applied to such measures.
4.1.1. Case I
Let $N\geq 1$. Suppose that $(a_{n}),(b_{n})$ are $N$-periodically modulated Jacobi parameters such that $|\operatorname{tr}\mathfrak{X}_{0}(0)|<2$ and for some $r\geq 1$
(4.5) $\bigg{(}\frac{a_{n-1}}{a_{n}}\bigg{)},\bigg{(}\frac{b_{n}}{a_{n}}\bigg{)},\bigg{(}\frac{1}{a_{n}}\bigg{)}\in\mathcal{D}_{r}^{N}.$
Then according to [35, Theorem A] the measure $\mu$ is determinate iff $\sum_{n=0}^{\infty}1/a_{n}=\infty$. If it is the case then by [35, Corollary 8] the measure $\mu$ is purely absolutely continuous on $\mathbb{R}$ with a continuous positive density. Thus, by Remark 2.5 the hypothesis (2.13) holds true for any $x\in\mathbb{R}$. Next, by [38, Theorem 1.2],
(4.6) $\lim_{n\to\infty}\frac{1}{\rho_{n}}K_{n}(x,x;\mu)=\frac{\upsilon(x)}{{\mu^{\prime}_{\mathrm{ac}}}(x)}$
locally uniformly with respect to $x\in\mathbb{R}$, where
(4.7) $\rho_{n}=\sum_{j=0}^{n-1}\frac{\alpha_{j}}{a_{j}}\quad\text{and}\quad\upsilon(x)\equiv\frac{1}{\pi N}\frac{|\operatorname{tr}\mathfrak{X}_{0}^{\prime}(0)|}{\sqrt{-\operatorname{discr}\mathfrak{X}_{0}(0)}}.$
Let $K\subset\mathbb{R}$ be compact. Then by (4.6) the limit (2.15) exists uniformly on $K$, and by Remark 2.6 the hypothesis (2.12) is satisfied for $K$. Therefore, by Theorem 2.4, Nevai’s condition holds uniformly on compact subsets of $\mathbb{R}$. Let us note that for $r=1$, the hypotheses (2.14) and (2.15) are also implied by [38, Theorem 1.3].
Let us comment some applications of our findings. In view of (4.6) the hypotheses of Corollary 3.1 are satisfied for any compact $K\subset\mathbb{R}$. Next, by [40, Theorem 7.1],
[…]
Therefore, the hypotheses of Theorem 3.2 are satisfied. In view of Remark 3.5 the hypotheses of Theorem 3.6 are satisfied. Moreover, Remark 3.8 implies that (3.24) holds provided
(4.9) $\lim_{n\to\infty}\frac{a_{n}}{n}=0.$
Let us remark that Hermite, Freud and Meixner–Pollaczek polynomials (see [32, Section 5] for details) belong to the case I and they satisfy (4.5).
4.1.2. Case III
Let $N\geq 1$. Suppose that $(a_{n}),(b_{n})$ are $N$-periodically modulated Jacobi parameters such that $|\operatorname{tr}\mathfrak{X}_{0}(0)|>2$, and for some $r\geq 1$,
(4.10) $\bigg{(}\frac{a_{n-1}}{a_{n}}\bigg{)},\bigg{(}\frac{b_{n}}{a_{n}}\bigg{)},\bigg{(}\frac{1}{a_{n}}\bigg{)}\in\mathcal{D}_{r}^{N}.$
Then according to [39, Theorem 5.3] (cf. [39, Remark 5.4]) the measure $\mu$ is determinate, the measure $\mu$ is purely discrete on $\mathbb{R}$ and $\operatornamewithlimits{supp}(\mu)$ has no finite accumulation points. Thus, in view of Theorem 2.3 Nevai’s condition holds uniformly on compact subsets of $\operatornamewithlimits{supp}(\mu)$.
Since ${\mu^{\prime}_{\mathrm{ac}}}(x)\equiv 0$ there seems to be no interesting applications of Section 3 for this case.
Let us remark that Meixner polynomials (see Example 4.2 for details) belong to the case III and they satisfy (4.10).
4.1.3. Case IIII(a)
Let $N\geq 1$. Suppose that $(a_{n}),(b_{n})$ are $N$-periodically modulated Jacobi parameters such that $|\operatorname{tr}\mathfrak{X}_{0}(0)|=2$, and $\mathfrak{X}_{0}(0)$ is diagonalizable. Then $\mathfrak{X}_{0}(0)=\varepsilon\operatorname{Id}$ for some $\varepsilon\in\{-1,1\}$. Suppose that
(4.11) $\bigg{(}a_{n}\Big{(}\frac{a_{n-1}}{a_{n}}-\frac{\alpha_{n-1}}{\alpha_{n}}\Big{)}\bigg{)},\bigg{(}a_{n}\Big{(}\frac{b_{n}}{a_{n}}-\frac{\beta_{n}}{\alpha_{n}}\Big{)}\bigg{)},\bigg{(}\frac{1}{a_{n}}\bigg{)}\in\mathcal{D}_{1}^{N}.$
Then according to [37, Proposition 4] the measure $\mu$ is determinate. Moreover, [33, Proposition 9] implies that
$h(x)=\lim_{j\to\infty}a_{jN+N-1}^{2}\operatorname{discr}X_{jN}(x)$
exists. In view of the proof of [33, Corollary 1], $h$ is a polynomial of degree $2$ with negative leading coefficient. Let $\Lambda_{-}=h^{-1}\big{(}(-\infty,0)\big{)}$ and $\Lambda_{+}=h^{-1}\big{(}(0,+\infty)\big{)}$. If
(4.12) $\lim_{n\to\infty}(a_{n+N}-a_{n})=0,$
then by [33, Theorem D] the measure $\mu$ is purely absolutely continuous on $\Lambda_{-}$ with positive continuous density. Thus, by Remark 2.5 the hypothesis (2.13) holds true for any $x\in\Lambda_{-}$. Next, by [37, Theorem 8 and Remark 1]
(4.13) $\lim_{n\to\infty}\frac{1}{\rho_{n}}K_{n}(x,x;\mu)=\frac{\upsilon(x)}{{\mu^{\prime}_{\mathrm{ac}}}(x)}$
locally uniformly with respect to $x\in\Lambda_{-}$, where
(4.14) $\rho_{n}=\sum_{j=0}^{n-1}\frac{\alpha_{j}}{a_{j}}\quad\text{and}\quad\upsilon(x)=\frac{1}{4\pi N\alpha_{N-1}}\frac{|h^{\prime}(x)|}{\sqrt{-h(x)}}\mathds{1}_{\Lambda_{-}}(x)$
(cf. [37, formula (7.12)]). Let $K\subset\Lambda_{-}$ be compact. Then by (4.13) the limit (2.15) exists uniformly on $K$. Thus, by Remark 2.6 the hypothesis (2.12) is satisfied for $K$. Therefore, by Theorem 2.4, Nevai’s condition holds uniformly on compact subsets of $\Lambda_{-}$. Let us note that the hypotheses (2.14) and (2.15) are also implied by [37, Remark 1 and Theorem D]. Next, by [39, Theorem C], the measure $\mu$ is purely discrete on $\Lambda_{+}$ and the only accumulation points of $\operatornamewithlimits{supp}(\mu)\cap\Lambda_{+}$ are contained in the set $\partial\Lambda_{-}=\{x:h(x)=0\}$ containing at most two points. Thus, by Theorem 2.3, Nevai’s condition holds uniformly on compact subsets of $\operatornamewithlimits{supp}(\mu)\setminus\partial\Lambda_{-}$.
Let us comment some applications of our findings. In view of (4.13) the hypotheses of Corollary 3.1 are satisfied for any compact $K\subset\Lambda_{-}$. Next, by [40, Theorem 7.2]
[…]
Therefore, the hypotheses of Theorem 3.2 are satisfied. In view of Remark 3.5 the hypotheses of Theorem 3.6 are satisfied. Moreover, Remark 3.8 implies that (3.24) is satified. Indeed, it is enough to prove that
(4.16) $\lim_{n\to\infty}\frac{a_{n}}{n}=0.$
Observe that for any $i\in\{0,1,\ldots,N-1\}$ we have by Stolz-Cesàro theorem and (4.12)
$\lim_{n\to\infty}\frac{a_{nN+i}}{nN+i}=\frac{1}{N}\lim_{n\to\infty}(a_{(n+1)N+i}-a_{nN+i})=0,$
which leads to (4.16).
Let us remark that generalized Hermite polynomials (see Example 4.3 for details) belong to the case IIII(a) and they satisfy (4.11) and (4.12).
4.1.4. Case IIII(b)
Let $N\geq 1$. Suppose that $(a_{n}),(b_{n})$ are $N$-periodically modulated Jacobi parameters such that $\mathfrak{X}_{0}(0)$ is non-diagonalizable. Let $\gamma=(\gamma_{n}:n\in\mathbb{N})$ be a sequence of positive numbers tending to infinity and satisfying
(4.17) $\bigg{(}\sqrt{\gamma_{n}}\Big{(}\sqrt{\frac{\alpha_{n-1}}{\alpha_{n}}}-\sqrt{\frac{\gamma_{n-1}}{\gamma_{n}}}\Big{)}\bigg{)},\bigg{(}\frac{1}{\sqrt{\gamma_{n}}}\bigg{)}\in\mathcal{D}_{1}^{N},$
and
(4.18) $\lim_{n\to\infty}\big{(}\sqrt{\gamma_{n+N}}-\sqrt{\gamma_{n}}\big{)}=0.$
Suppose that
(4.19) $\bigg{(}\sqrt{\gamma_{n}}\Big{(}\frac{\alpha_{n-1}}{\alpha_{n}}-\frac{a_{n-1}}{a_{n}}\Big{)}\bigg{)},\bigg{(}\sqrt{\gamma_{n}}\Big{(}\frac{\beta_{n}}{\alpha_{n}}-\frac{b_{n}}{a_{n}}\Big{)}\bigg{)},\bigg{(}\frac{\gamma_{n}}{a_{n}}\bigg{)}\in\mathcal{D}_{1}^{N}$
and for $\varepsilon=\operatorname{sign}({\operatorname{tr}\mathfrak{X}_{0}(0)})$
(4.20) $\bigg{(}\gamma_{n}\big{(}1-\varepsilon[\mathfrak{X}_{n}(0)]_{11}\big{)}\Big{(}\frac{\alpha_{n-1}}{\alpha_{n}}-\frac{a_{n-1}}{a_{n}}\Big{)}-\gamma_{n}\varepsilon\Big{(}\frac{\beta_{n}}{\alpha_{n}}-\frac{b_{n}}{a_{n}}\Big{)}\bigg{)}\in\mathcal{D}_{1}^{N}.$
Then according to [41, Theorem A]
$h(x)=\lim_{j\to\infty}\gamma_{jN+N-1}\operatorname{discr}X_{jN}(x)$
exists and is a polynomial of degree at most one. Let $\Lambda_{-}=h^{-1}\big{(}(-\infty,0)\big{)}$ and $\Lambda_{+}=h^{-1}\big{(}(0,+\infty)\big{)}$. Suppose that $\Lambda_{-}\cup\Lambda_{+}\neq\emptyset$ and that the measure $\mu$ is determinate (for detailed discussion of the second property see [41, Section 9]).
If $\Lambda_{-}\neq\emptyset$ we define
(4.21) $\rho_{n}=\sum_{j=0}^{n-1}\frac{\sqrt{\alpha_{j}\gamma_{j}}}{a_{j}}\quad\text{and}\quad\upsilon(x)=\frac{\sqrt{\alpha_{N-1}}}{\pi N}\frac{|\operatorname{tr}\mathfrak{X}_{0}^{\prime}(0)|}{\sqrt{-h(x)}}\mathds{1}_{\Lambda_{-}}(x).$
Then according to [41, Theorem B] the measure $\mu$ is determinate iff $\rho_{n}\to\infty$. Hence, according to [40, Corollary 7.4], the measure $\mu$ is purely absolutely continuous on $\Lambda_{-}$ with continuous positive density and
(4.22) $\lim_{n\to\infty}\frac{1}{\rho_{n}}K_{n}(x,x;\mu)=\frac{\upsilon(x)}{{\mu^{\prime}_{\mathrm{ac}}}(x)}$
locally uniformly with respect to $x\in\Lambda_{-}$. Thus, by Remark 2.5 the hypothesis (2.13) holds true for any $x\in\Lambda_{-}$. Let $K\subset\Lambda_{-}$ be compact. Then by (4.22) the limit (2.15) exists uniformly on $K$. Let $K\subset\Lambda_{-}$ be compact. Consequently, by Theorem 2.4, Nevai’s condition holds uniformly on compact subsets of $\Lambda_{-}$. Let us note that the hypotheses (2.14) and (2.15) under stronger assumptions follows from [36, Theorem D]. Next, if $\Lambda_{+}\neq\emptyset$, then by [41, Theorem A] the measure $\mu$ is purely discrete on $\Lambda_{+}$ and the only accumulation points of $\operatornamewithlimits{supp}(\mu)\cap\Lambda_{+}$ are contained in the set $\partial\Lambda_{-}=\{x:h(x)=0\}$ containing at most one point. Thus, by Theorem 2.3, Nevai’s condition holds uniformly on compact subsets of $\operatornamewithlimits{supp}(\mu)\setminus\partial\Lambda_{-}$.
Let us comment some applications of our findings. In view of (4.22) the hypotheses of Corollary 3.1 are satisfied for any compact $K\subset\Lambda_{-}$. Next, by [40, Theorem 7.3],
[…]
Therefore, the hypotheses of Theorem 3.2 are satisfied. In view of Remark 3.5 the hypotheses of Theorem 3.6 are satisfied. Moreover, Remark 3.8 implies that (3.24) holds provided
(4.24) $\lim_{n\to\infty}\frac{a_{n}}{n\sqrt{\gamma_{n}}}=0.$
Let us remark that Laguerre-type polynomials (see Example 4.4 for details) belong to the case IIII(b) and they satisfy (4.17), (4.18), (4.19), (4.20) for $\gamma_{n}=a_{n}$.
4.2. Periodic blend
We say that Jacobi parameters $(a_{n}),(b_{n})$ are $N$-periodically blended, if there exist $N$-periodic sequences $(\alpha_{n}),(\beta_{n})$ of positive and real numbers, respectively, such that
$\lim_{j\to\infty}|a_{j(N+2)+i}-\alpha_{i}|=0,\quad\lim_{j\to\infty}|b_{j(N+2)+i}-\beta_{i}|=0,\quad i=0,1,\ldots,N-1,$
and $b_{j(N+2)+N}=b_{j(N+2)+N+1}=0$, and
(4.25) $\lim_{j\to\infty}a_{j(N+2)+N}=\infty,\quad\lim_{j\to\infty}\frac{a_{j(N+2)+N}}{a_{j(N+2)+N+1}}=1.$
Next, we define $(N+2)$-step transfer matrix by
$X_{n}(x)=B_{N+1+n}(x)\ldots B_{n+1}(x)B_{n}(x),\quad n\geq 1.$
Then by [38, Section 3.3] we have
$\mathcal{X}_{1}(x)=\lim_{j\to\infty}X_{j(N+2)+1}(x)=\begin{pmatrix}0&-1\\ -\frac{\alpha_{N-1}}{\alpha_{0}}&-\frac{2x-\beta_{0}}{\alpha_{0}}\end{pmatrix}\mathfrak{B}_{N-1}(x)\ldots\mathfrak{B}_{2}(x)\mathfrak{B}_{1}(x).$
Let $h(x)=\operatorname{discr}\mathcal{X}_{1}(x)$ and set
$\Lambda_{-}=h^{-1}\big{(}(-\infty,0)\big{)}\quad\text{and}\quad\Lambda_{+}=h^{-1}\big{(}(0,+\infty)\big{)}.$
By [38, Theorem 3.13]
$\Lambda_{-}=\bigcup_{j=1}^{N}I_{j},$
where $I_{j}$ are disjoint open non-empty bounded intervals whose closures might touch each other.
Notice that by (4.25) the condition (1.4) is violated. Therefore, measures $\mu$ corresponding to $N$-periodically blended Jacobi parameters have unbounded supports. This class has been introduced in [1] as an explicit example of measures such that $\operatornamewithlimits{supp}({\mu_{\mathrm{ac}}})$ is a closure of $\Lambda_{-}$. Later, this class has been studied in more detail in [39, 38, 35].
Suppose that $(a_{n}),(b_{n})$ are $N$-periodically blended Jacobi parameters such that for some $r\geq 1$
$\bigg{(}\frac{1}{a_{n}}\bigg{)},\bigg{(}\frac{b_{n}}{a_{n}}\bigg{)}\in\mathcal{D}_{r}^{N+2}\quad\text{and}\quad\bigg{(}\frac{a_{n(N+2)+N}}{a_{n(N+2)+N+1}}\bigg{)}\in\mathcal{D}_{r}^{1}.$
Then according to [35, Corollary 9] the moment problem for $\mu$ is determinate and the measure $\mu$ is purely absolutely continuous on $\Lambda_{-}$ with continuous positive density. Thus, by Remark 2.5 the hypothesis (2.13) holds true for any $x\in\Lambda_{-}$. Next, by [38, Theorem 4.10] (cf. Theorem 3.13, Corollary 3.15 and Remark 4.14 from [38])
(4.26) $\lim_{n\to\infty}\frac{1}{\rho_{n}}K_{n}(x,x;\mu)=\frac{\upsilon(x)}{{\mu^{\prime}_{\mathrm{ac}}}(x)}$
locally uniformly with respect to $x\in\Lambda_{-}$, where
(4.27) $\rho_{n}=n\quad\text{and}\quad\upsilon(x)=\frac{1}{\pi(N+2)}\frac{|\operatorname{tr}\mathcal{X}_{1}^{\prime}(0)|}{\sqrt{-\operatorname{discr}\mathcal{X}_{1}(0)}}\mathds{1}_{\Lambda_{-}}(x).$
Thus, by Remark 2.6 the hypothesis (2.12) is satisfied for any compact $K\subset\Lambda_{-}$. Therefore, by Theorem 2.4, Nevai’s condition holds uniformly on compact subsets of $\Lambda_{-}$. Let us note that for $r=1$, the hypotheses (2.14) and (2.15) are also implied by [38, Theorem 5.17]. Next, by [39, Theorem 5.1] (cf. [39, Remark 5.2]), the measure $\mu$ is purely discrete on $\Lambda_{+}$ and the only accumulation points of $\operatornamewithlimits{supp}(\mu)\cap\Lambda_{+}$ are contained in the set $\partial\Lambda_{-}=\{x:h(x)=0\}$ containing at most $2N$ points. Thus, by Theorem 2.3, Nevai’s condition holds uniformly on compact subsets of $\operatornamewithlimits{supp}(\mu)\setminus\partial\Lambda_{-}$.
In view of (4.26) the hypotheses of Corollary 3.1 are satisfied for any compact $K\subset\Lambda_{-}$. Let us remark that weak convergence of the Christoffel–Darboux kernels for periodic blend has not been studied yet, but is seems that the techniques of [40] can be applied here.
4.3. Examples of explicit measures
In this section we present a few examples of well-known measures with periodically modulated Jacobi parameters satisfying the appropriate regularity conditions from Section 4.1.
Example 4.1 (Freud weights).
For $\gamma\geq 1$ consider probability measure
${\>\rm d}\mu=c_{\gamma}\textrm{e}^{-|x|^{\gamma}}{\>\rm d}\lambda(x),$
where $c_{\gamma}$ is the normalizing constant. Notice that for $\gamma=2$ we recover a normal distribution. According to [32, Example 3] the Jacobi parameters of the determinate measure $\mu$ satisfy (4.3) with $N=1$ and $\alpha_{n}\equiv 1,\beta_{n}\equiv 0$. Then by (4.4) we have
(4.28) $\mathfrak{X}_{0}(x)=\begin{pmatrix}0&1\\ -1&x\end{pmatrix}.$
Thus, $\operatorname{tr}\mathfrak{X}_{0}(0)=0$. Therefore, we are in the setup of Section 4.1.1. Moreover, the condition (4.5) is satisfied for $r=1$. Let $\rho_{n}$ and $\upsilon$ be defined in (4.21). By using (4.28) we immediately get $\upsilon(x)\equiv\frac{1}{2\pi}$. By Stolz–Cesàro Theorem and [32, Example 3] one can check that
(4.29) $\lim_{n\to\infty}\frac{\rho_{n}}{\tilde{\rho}_{n}}=1\quad\text{where}\quad\tilde{\rho}_{n}=\frac{1}{2}\Bigg{(}\frac{\Gamma\big{(}\tfrac{\gamma}{2}\big{)}\Gamma\big{(}\tfrac{1}{2}\big{)}}{\Gamma\big{(}\tfrac{1+\gamma}{2}\big{)}}\Bigg{)}^{\tfrac{1}{\gamma}}\sum_{j=1}^{n}j^{-\tfrac{1}{\gamma}}.$
Moreover, by [32, Example 3] the condition (4.9) is satisfied iff $\gamma>1$.
Example 4.2 (Meixner polynomials).
For any $s>0$ and $p\in(0,1)$ the negative binomial distribution is the probability measure
$\mu=(1-p)^{s}\sum_{k=0}^{\infty}\binom{k+s-1}{k}p^{k}\delta_{k}.$
Notice that for $s=1$ we recover geometric distributions. The orthonormal polynomials in $L^{2}(\mu)$ are called Meixner polynomials and they satisfy
$a_{n}=\frac{\sqrt{(n+1)(n+s)p}}{1-p},\quad b_{n}=\frac{n+(n+s)p}{1-p},$
see e.g. [12, formula (9.10.4)]. It is easy to check that (4.3) is satisfied for $N=1$ and $\alpha_{n}\equiv 1,\beta_{n}\equiv\sqrt{p}+\frac{1}{\sqrt{p}}>2$. Then by (4.4) we have
$\mathfrak{X}_{0}(x)=\begin{pmatrix}0&1\\ -1&x-\sqrt{p}-\frac{1}{\sqrt{p}}\end{pmatrix}.$
Thus, we are in the setup of Section 4.1.2. Moreover, it is easy to show that (4.10) is satisfied for $r=1$.
Example 4.3 (Generalized Hermite polynomials).
For any $t>-1$ consider probability measure
${\>\rm d}\mu=\frac{1}{\Gamma(\tfrac{1+t}{2})}|x|^{t}\textrm{e}^{-x^{2}}{\>\rm d}\lambda(x).$
Notice that for $t=0$ we recover a normal distribution. According to [32, Example 1] the Jacobi parameters of $\mu$ satisfy (4.3) with $N=1$ and $\alpha_{n}\equiv 1,\beta_{n}\equiv 0$. However, unless $t=0$ the condition (4.10) is not satisfied for $N=1$. Again by [32, Example 1], the condition (4.11) is satisfied for $N=2$ and we are in the setup of Section 4.1.3. Let $\rho_{n}$ and $\upsilon$ be defined in (4.14). By Stolz–Cesàro Theorem one can check that
$\lim_{n\to\infty}\frac{\rho_{n}}{\tilde{\rho}_{n}}=1\quad\text{where}\quad\tilde{\rho}_{n}=\sqrt{2}\sum_{j=1}^{n}j^{-\tfrac{1}{2}}$
Moreover, [32, Corollary 3] (see also [37, Theorem 6]) implies that $\upsilon(x)=\frac{1}{2\pi}\mathds{1}_{\mathbb{R}\setminus\{0\}}(x)$. Finally, the condition (4.16) is satisfied.
Example 4.4 (Laguerre-type polynomials).
For any $\gamma>-1$ and $\kappa\in\{2,3,\ldots\}$ consider the probability measure
${\>\rm d}\mu=c_{\gamma,\kappa}x^{\gamma}\textrm{e}^{-x^{\kappa}}\mathds{1}_{(0,\infty)}(x){\>\rm d}\lambda(x),$
where $c_{\gamma,\kappa}$ is the normalizing constant. According to [36, Section 10.1.1] the Jacobi parameters of the measure $\mu$ satisfy (4.3) with $N=1$ and $\alpha_{n}\equiv 1,\beta_{n}\equiv 2$. Then by (4.4) we have
$\mathfrak{X}_{0}(x)=\begin{pmatrix}0&1\\ -1&x-2\end{pmatrix}.$
Hence, we are in the setup of Section 4.1.4. Moreover, again by [36, Section 10.1.1], the hypotheses (4.17), (4.18), (4.19) and (4.20) are satisfied for $\gamma_{n}=a_{n}$. Let $\rho_{n}$ and $\upsilon$ be defined in (4.21). By Stolz–Cesàro Theorem and [36, Section 10.1.1] one can check that
$\lim_{n\to\infty}\frac{\rho_{n}}{\tilde{\rho}_{n}}=1\quad\text{where}\quad\tilde{\rho}_{n}=\frac{1}{2}\bigg{(}\frac{2(2\kappa)!!}{\kappa(2\kappa-1)!!}\bigg{)}^{\tfrac{1}{2\kappa}}\sum_{j=1}^{n}j^{-\tfrac{1}{2\kappa}}.$
Moreover, by [36] (see Section 10.1.1, Corollary 3.4 and formula (3.1)) we have $h(x)=-4x$, from which one can verify that $\upsilon(x)=\frac{1}{2\pi}\frac{1}{\sqrt{x}}\mathds{1}_{(0,\infty)}(x)$. Finally, the condition (4.24) is satisfied.